OpenStudy (anonymous):

A baseball is thrown up into the air. The ball's height in feet is modeled by the equation h(t) = -16t2 + 32t + 10. (h = height, t = time in seconds) Find the maximum height of the ball. a) 10 feet b) 23 feet c) 2.3 feet d) 26 feet

OpenStudy (anonymous):

The answer id D 26 feet.

OpenStudy (anonymous):

\[h(t) = -16t^2+32t+10\]\[dh/dt=-32t+32\]now make that = 0 \[-32t+32=0 \rightarrow t=1\]when t=1\[h(1)=-16(1)^2+32(1)+10 = 26\] so answer is D

OpenStudy (anonymous):

thanks